OpenStudy (anonymous):

I will give medals Which points are important when graphing sine and cosine functions?

OpenStudy (anonymous):

The "important" points: -The zeros, maximum and minimum points. The Sine curve has zeros at the beginning, middle and end of a cycle. The maximum happens at the 1/4 mark and the minimum appears at the 3/4 mark. -The Cosine curve begins and ends with the maximum. It has a minimum at the middle point. Zeros appear at the 1/4 and 3/4 mark of the cycle. http://home.windstream.net/okrebs/page73.html

OpenStudy (anonymous):

What I would do is this: when \[\theta=0\]cos(0)=1 sin(0)=0 ------------------------------------------\[\theta=\frac{ \Pi }{ 2 }\]cos(pi/2)=0 sin(pi/2)=1 ------------------------------------------\[\theta=3\Pi\]cos(3pi)=-1 sin(3pi)=0 ------------------------------------------\[\theta=\frac{ 3\Pi }{ 2 }\]cos(3pi/2)=0 sin (3pi/2)=-1 ------------------------------------------\[\theta=2\Pi\] cos(2pi)=1 sin(2pi)=0 Remember that one cycle depends from 0 to 2pi. The equivalents are just from a unit circle which is from above. so the coordinate system should look like this Based on the coordinate system, Just plot which function you're going to take first, sine function or cosine function